value of cot 1358 degree?
Answer by stanbon(75887)

value of cot 1358 degree?
-----
1358 = 3*360 + 278
----
cot(1358) = cot(278) = 1/tan(278) = -0.1405
======
Cheers,
Stan H.
=============